Use the price-demand equation to determine whether demand is elastic, is inelastic, or has unit elasticity at the Indicated v Need help with 13,14,15
Thank you :)
Will upvote!
Answer #1

Elasticity of demand. E (P) - - Pf(P) f(P 298–4x52 Elastic (3 x = f(P) = 298 - 4P ; P = 52. f(P) = -4 F(P) = _ (S2)(-4) be